OpenStudy (anonymous):

The coefficent of static friction for steel on ice is 0.1. the force needed to set 70 kg skater in motion is about? @incredebility

OpenStudy (anonymous):

static friction= 0.1 *W

OpenStudy (anonymous):

w= mg..so g= 9.8 m/s2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The minimum force required to put an object into motion has to be greater than or equal to the frictional force (Ff) and this is given by Ff=$R, where $ is friction, and R is normal contact force or mg or 9.81*70= 686.7Newtons. Then Ff or minimum force= 0.1*686.7= 68.67N.